# Wilsonian Renormalization of Neural Network Gaussian Processes

### v1

@@ 5 Discussion
In this work, we applied the powerful formalism of Wilsonian RG to Gaussian Process regression. In particular, treating the output of a trained DNN as a statistical ensemble, we split the associated fluctuations into high- and low-energy modes (based on their stiffness in a random network). Provided the variance of the higher modes is much smaller than the ridge parameter, these may be integrated out in a tractable manner leading to a renormalized GP regression problem. In simple settings where the features associated with those modes are Gaussian, this results in a renormalized ridge parameter. In more general settings, one obtains an effective GP regression problem where the ridge parameter becomes spatially dependent. We then applied this to a simple toy model, in which the effects determined via our RG approach are demonstrably missed in the standard eigen-learning framework canatar .
@@
It would be interesting to generalize our formalism to include fluctuation effects over datasets by keeping track of $y_{>}^{2}$ dependencies. Furthermore, it would be interesting to extend our one-step RG treatment for the non-Gaussian case into a fully-fledged functional RG flow. This may enable us to define a notion of relevancy and irrelevancy applicable to generic non-Gaussian features and, potentially, a notion of universality classes for DNNs. Similarly, our formalism could be extended to finite neural networks where additional interaction terms would appear corresponding to finite-width corrections to the Gaussian prior, and investigate universality in those corrections.
Acknowledgements
@@
JNH was supported by the National Science Foundation under Grant No. NSF PHY-1748958 and by the Gordon and Betty Moore Foundation through Grant No. GBMF7392. AM acknowledges support from Perimeter Institute, which is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Colleges and Universities. ZR acknowledges support from ISF Grant 2250/19. JNH, RJ, and AM wish to thank the Mainz Institute for Theoretical Physics for hosting the 2022 workshop “A Deep-Learning Era of Particle Theory”, where this collaboration was initiated, as well as the Aspen Center for Physics (ACP, supported by National Science Foundation grant PHY-2210452) for hospitality during the 2023 winter conference “Theoretical Physics for Machine Learning”. JNH, and ZR acknowledge further support from ACP through the 2023 summer program “Theoretical Physics for Deep Learning”. This research was also supported in part by grants NSF PHY-1748958 and PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP), through the participation of AM, JNH, and ZR in the Fall 2023 program “Deep Learning from the Perspective of Physics and Neuroscience.”
@@ Appendix A Spectral decomposition of the Gaussian action
$S=\sum_{m=1}^{M}\frac{1}{2}\int\!\mathrm{d}\mu_{x}\mathrm{d}\mu_{x^{\prime}}\,f_{m}(x)K^{-1}(x,x^{\prime})f_{m}(x^{\prime})+\frac{\eta}{\sigma^{2}}\int\!\mathrm{d}\mu_{x}\sum_{m}\frac{\left(f_{m}(x)-y(x)\right)^{2}}{2}\leavevmode\nobreak\ .$ (64)
@@
$=\int\!\mathrm{d}\mu_{x}\mathrm{d}\mu_{x^{\prime}}\left(\sum_{k}f_{mk}\phi_{k}(x)\right)K^{-1}(x,x^{\prime})\left(\sum_{k^{\prime}}f_{mk^{\prime}}\phi_{k^{\prime}}(x^{\prime})\right)$
@@
$\int\!\mathrm{d}\mu_{x}\left(f_{m}(x)-y(x)\right)^{2}$ $=\int\!\mathrm{d}\mu_{x}\left[\sum_{k}\left(f_{mk}-y_{k}\right)\phi_{k}(x)\right]\left[\sum_{k^{\prime}}\left(f_{mk^{\prime}}-y_{k^{\prime}}\right)\phi_{k^{\prime}}(x)\right]$ (66)
@@
$\int$ $\mathrm{d}\mu_{x}\left(\sum_{k\leq\kappa}f_{mk}\phi_{k}+\sum_{k>\kappa}f_{mk}\phi_{k}\right)\left(\sum_{k^{\prime}\leq\kappa}\lambda_{k^{\prime}}^{-1}f_{mk^{\prime}}\phi_{k^{\prime}}+\sum_{k^{\prime}>\kappa}\lambda_{k^{\prime}}^{-1}f_{mk^{\prime}}\phi_{k}^{\prime}\right)$ (67)
$=\left(\sum_{k,k^{\prime}\leq\kappa}+\sum_{k,k^{\prime}>\kappa}+\sum_{\begin{subarray}{c}k\leq\kappa\\ k^{\prime}>\kappa\end{subarray}}+\sum_{\begin{subarray}{c}k>\kappa\\ k^{\prime}\leq\kappa\end{subarray}}\right)\lambda_{k^{\prime}}^{-1}f_{mk}f_{mk^{\prime}}\int\!\mathrm{d}\mu_{x}\,\phi_{k}(x)\phi_{k^{\prime}}(x)$

### v2

@@ V Discussion
In this work, we applied the powerful formalism of Wilsonian RG to Gaussian Process regression. In particular, treating the output of a trained DNN as a statistical ensemble, we split the associated fluctuations into high- and low-energy modes (based on their stiffness in a random network). Provided the variance of the higher modes is much smaller than the ridge parameter, these may be integrated out in a tractable manner leading to a renormalized GP regression problem. In simple settings where the features associated with those modes are Gaussian, this results in a renormalized ridge parameter. In more general settings, one obtains an effective GP regression problem where the ridge parameter becomes spatially dependent. We then applied this to a simple toy model, in which the effects determined via our RG approach are demonstrably missed in the standard eigen-learning framework [5].
@@
It would be interesting to generalize our formalism to include higher order contributions in the Non-Gaussianity as well as fluctuation effects over datasets by keeping track of $y_{>}^{2}$ dependencies. Such an effort would also benefit from diagrammatic reformulation of our derivation. Such a formalism may enable us to define a notion of relevancy and irrelevancy applicable to generic non-Gaussian features and, potentially, a notion of universality classes for DNNs. Similarly, our formalism could be extended to finite neural networks where additional interaction terms would appear corresponding to finite-width corrections to the Gaussian prior, and investigate universality in those corrections.
Acknowledgements.
@@
We wish to thank Marc Klinger and Alexander Stapleton for useful discussions regarding the Bayesian Renormalization Group and related topics. JNH was supported by the National Science Foundation under Grant No. NSF PHY-1748958 and by the Gordon and Betty Moore Foundation through Grant No. GBMF7392. AM acknowledges support from Perimeter Institute, which is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Colleges and Universities. ZR acknowledges support from ISF Grant 2250/19. JNH, RJ, and AM wish to thank the Mainz Institute for Theoretical Physics for hosting the 2022 workshop “A Deep-Learning Era of Particle Theory”, where this collaboration was initiated, as well as the Aspen Center for Physics (ACP, supported by National Science Foundation grant PHY-2210452) for hospitality during the 2023 winter conference “Theoretical Physics for Machine Learning”. JNH, and ZR acknowledge further support from ACP through the 2023 summer program “Theoretical Physics for Deep Learning”. This research was also supported in part by grants NSF PHY-1748958 and PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP), through the participation of AM, JNH, and ZR in the Fall 2023 program “Deep Learning from the Perspective of Physics and Neuroscience.”
@@ Appendix A Spectral decomposition of the Gaussian action
$S=$ $\sum_{m=1}^{M}\frac{1}{2}\int\!\mathrm{d}\mu_{x}\mathrm{d}\mu_{x^{\prime}}\,f_{m}(x)K^{-1}(x,x^{\prime})f_{m}(x^{\prime})$ (81)
$+\frac{\eta}{\sigma^{2}}\int\!\mathrm{d}\mu_{x}\sum_{m}\frac{\left(f_{m}(x)-y(x)\right)^{2}}{2}\leavevmode\nobreak\ .$
@@
$=\int\!\mathrm{d}\mu_{x}\mathrm{d}\mu_{x^{\prime}}$
×$\qquad\times\left(\sum_{k}f_{mk}\phi_{k}(x)\right)K^{-1}(x,x^{\prime})\left(\sum_{k^{\prime}}f_{mk^{\prime}}\phi_{k^{\prime}}(x^{\prime})\right)$
@@
$\int$ $\mathrm{d}\mu_{x}\left(f_{m}(x)-y(x)\right)^{2}$ (83)
$=\int\!\mathrm{d}\mu_{x}\left[\sum_{k}\left(f_{mk}-y_{k}\right)\phi_{k}(x)\right]$
×$\qquad\quad\times\left[\sum_{k^{\prime}}\left(f_{mk^{\prime}}-y_{k^{\prime}}\right)\phi_{k^{\prime}}(x)\right]$
@@
$\int$ $\mathrm{d}\mu_{x}\left(\sum_{k\leq\kappa}f_{mk}\phi_{k}+\sum_{k>\kappa}f_{mk}\phi_{k}\right)$ (84)
×$\qquad\times\left(\sum_{k^{\prime}\leq\kappa}\lambda_{k^{\prime}}^{-1}f_{mk^{\prime}}\phi_{k^{\prime}}+\sum_{k^{\prime}>\kappa}\lambda_{k^{\prime}}^{-1}f_{mk^{\prime}}\phi_{k}^{\prime}\right)$
$=\left(\sum_{k,k^{\prime}\leq\kappa}+\sum_{k,k^{\prime}>\kappa}+\sum_{\begin{subarray}{c}k\leq\kappa\\ k^{\prime}>\kappa\end{subarray}}+\sum_{\begin{subarray}{c}k>\kappa\\ k^{\prime}\leq\kappa\end{subarray}}\right)\lambda_{k^{\prime}}^{-1}f_{mk}f_{mk^{\prime}}$
×$\qquad\qquad\qquad\qquad\qquad\qquad\times\int\!\mathrm{d}\mu_{x}\,\phi_{k}(x)\phi_{k^{\prime}}(x)$
@@ Appendix B Simplifying interactions in the replica limit
Appendix C $\delta W_{>}(x)$ dependent terms
To evaluate (66), it is convenient to split $\delta W_{>}$ into components involving a single greater feature mode, which we denote $\delta W_{>}^{(1)}$, and those involving two greater feature modes, which we denote $\delta W_{>}^{(2)}$, cf. (63).
Now, consider first the term in (66), proportional to $\frac{\delta W^{(2)}_{>}}{\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}$. As it is odd in the greater feature mode, we have
-- $\int\!\mathcal{D}\varphi_{>}{\cal N}[0,1;\varphi_{>}]\left(1-\frac{1}{\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}+\frac{1}{2}\left[\frac{1}{\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}\right]^{2}\right)$ (87)
×$\qquad\qquad\times\frac{\delta W^{(2)}_{>}}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}$
$=\int\!\mathcal{D}\varphi_{>}{\cal N}[0,1;\varphi_{>}]\frac{1}{2\sigma^{4}}\left(\Phi_{<}^{\top}\Phi_{>}\right)^{2}\delta W^{(2)}_{>}$
$=\frac{1}{2\sigma^{4}}\left(\Phi_{<}^{\top}\bm{f}_{q^{\prime}}\right)^{2}\int\!\mathcal{D}\varphi_{>}{\cal N}[0,1;\varphi_{>}]\delta W^{(2)}_{>}\phi^{2}_{q^{\prime}}$
$=\frac{1}{\sigma^{4}}\left(\Phi_{<}^{\top}\bm{f}_{q^{\prime}}\right)^{2}W_{q^{\prime}q^{\prime}}\eqqcolon T_{1}\leavevmode\nobreak\ .$
As we are keeping only leading order in $W$, under the GP integration this term would simply be replaced by its average under the free greater mode action where it would yield the following additional contribution to the renormalized GP action
$C_{1}\coloneqq$ $\int\!\mathcal{D}\varphi_{<}\,P[\varphi_{<}]\exp\!\left\{-\frac{1+\delta W_{<}}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{<}\right\}$ (88)
×$\qquad\qquad\times\frac{1}{\sigma^{2}}\Phi_{<}^{\top}\Phi_{<}\frac{W_{q^{\prime}q^{\prime}}\lambda_{q^{\prime}}}{\sigma^{2}}\leavevmode\nobreak\ .$
Next we consider the term proportional to $\frac{\delta W^{(1)}_{>}}{\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}$. As it is even in the greater feature modes, we have
-- $\int\!\mathcal{D}\varphi_{>}{\cal N}[0,1;\varphi_{>}]\left(1-\frac{1}{\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}+\frac{1}{2}\left[\frac{1}{\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}\right]^{2}\right)$ (89)
×$\qquad\qquad\times\frac{\delta W^{(1)}_{>}}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}$
$=-\int\!\mathcal{D}\varphi_{>}{\cal N}[0,1;\varphi_{>}]\left(1+\frac{1}{2}\left[\frac{1}{\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}\right]^{2}\right)$
×$\qquad\qquad\times\frac{\delta W^{(1)}_{>}}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}$
$=-\frac{1}{2\sigma^{2}}\Phi_{<}^{\top}\bm{f}_{q^{\prime}}\int\!\mathcal{D}\varphi_{>}{\cal N}[0,1;\varphi_{>}]\delta W^{(1)}_{>}\varphi_{q^{\prime}}$
$\qquad\qquad-2\left[\frac{1}{2\sigma^{2}}\Phi_{<}^{\top}\bm{f}_{q^{\prime}}\right]^{3}\int\!\mathcal{D}\varphi_{>}{\cal N}[0,1;\varphi_{>}]\delta W^{(1)}_{>}\varphi_{q^{\prime}}^{3}$
$=-\frac{1}{2\sigma^{2}}\Phi_{<}^{\top}\bm{f}_{q^{\prime}}\left(\sum_{k_{1}\neq q^{\prime}}W_{k_{1},q^{\prime}}\varphi_{k_{1}}\right)$
$\qquad\qquad-2\left[\frac{1}{2\sigma^{2}}\Phi_{<}^{\top}\bm{f}_{q^{\prime}}\right]^{3}\left(3\sum_{k_{1}\neq q^{\prime}}W_{k_{1},q^{\prime}}\varphi_{k_{1}}\right)$
$\eqqcolon T_{2}+T_{3}\leavevmode\nobreak\ .$
Conveniently, the corresponding contribution ($C_{2}$, $C_{3}$, cf. (88)) to the renormalized GP action would be proportional to the average of these two terms under the free greater GP mode action, which vanishes due to their odd order in $\bm{f}_{q^{\prime}}$.
We next consider the term proportional to $\frac{\delta W^{(1)}_{>}}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{<}$ in (66). Using the fact that this is odd in the greater feature mode we have
$-\int\!\mathcal{D}\varphi_{>}{\cal N}[0,1;\varphi_{>}]\left(1-\frac{1}{\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}+\frac{1}{2}\left[\frac{1}{\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}\right]^{2}\right)$ (90)
×$\qquad\qquad\times\frac{\delta W^{(1)}_{>}}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{<}$
$=\int\!\mathcal{D}\varphi_{>}{\cal N}[0,1;\varphi_{>}]\frac{1}{\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}\frac{\delta W^{(1)}_{>}}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{<}$
$=\frac{1}{\sigma^{2}}\Phi_{<}^{\top}\bm{f}_{q^{\prime}}\frac{1}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{<}\int\!\mathcal{D}\varphi_{>}{\cal N}[0,1;\varphi_{>}]\delta W^{(1)}_{>}\varphi_{q^{\prime}}$
$\eqqcolon T_{4}\leavevmode\nobreak\ .$
However, the associated contribution ($C_{4}$) to the renormalized GP action again vanishes due to the odd order in $\bm{f}_{q^{\prime}}$.
Finally, we consider the term proportional to $\frac{\delta W^{(2)}_{>}}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{<}$, which is even in the greater feature mode, hence
$-\int\!\mathcal{D}\varphi_{>}{\cal N}[0,1;\varphi_{>}]\left(1-\frac{1}{\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}+\frac{1}{2}\left[\frac{1}{\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}\right]^{2}\right)$ (91)
×$\qquad\qquad\times\frac{\delta W^{(2)}_{>}}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{<}$
$=-\int\!\mathcal{D}\varphi_{>}{\cal N}[0,1;\varphi_{>}]\frac{1}{4\sigma^{2}}\left[\frac{1}{\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}\right]^{2}\delta W^{(2)}_{>}\Phi_{<}^{\top}\Phi_{<}$
$=-\frac{1}{4\sigma^{2}}\left[\frac{1}{\sigma^{2}}\Phi_{<}^{\top}\bm{f}_{q^{\prime}}\right]^{2}\Phi_{<}^{\top}\Phi_{<}\int\!\mathcal{D}\varphi_{>}{\cal N}[0,1;\varphi_{>}]\delta W^{(2)}_{>}\varphi_{q^{\prime}}^{2}$
$=-\frac{1}{2\sigma^{2}}\left[\frac{1}{\sigma^{2}}\Phi_{<}^{\top}\bm{f}_{q^{\prime}}\right]^{2}\Phi_{<}^{\top}\Phi_{<}W_{q^{\prime}q^{\prime}}\eqqcolon T_{5}\leavevmode\nobreak\ ,$
where the first term in the first line vanishes since $He_{2}(\phi_{q}^{2})$ integrates to zero under the standard normal measure. This yeilds the following contribution to the renormalized GP action
$C_{5}\coloneqq$ $-\int\!\mathcal{D}\varphi_{<}\,P[\varphi_{<}]\exp\!\left\{-\frac{1+\delta W_{<}}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{<}\right\}$ (92)
×$\qquad\qquad\times\frac{1}{2}\left[\frac{1}{\sigma^{2}}\Phi_{<}^{\top}\Phi_{<}\right]^{2}\frac{W_{q^{\prime}q^{\prime}}\lambda_{q^{\prime}}}{\sigma^{2}}\leavevmode\nobreak\ .$
However, both non-zero contributions ($C_{1}$, $C_{5}$) are proportional to $W\lambda_{q^{\prime}}/\sigma^{2}$, and hence contribute at order $[\delta c/\sigma_{c}^{2}]^{2}$. Integrating out those modes up to a total variance of ${\sum^{k_{min}}_{k_{max}}\delta_{c}=c}$ following the renormalization flow (70), $W_{kk^{\prime}}$ may be $O(c)$ and non-negligible; however, any contribution of the type $W_{kk^{\prime}}\delta c$ would still need to vanish in the Wilsonian, $\delta c\rightarrow 0$, limit.